# A Review of Graph-Powered Data Quality Applications for IoT Monitoring Sensor Networks

### v1

@@ 4.3.4 Clustering and Data Aggregation
In the context of data aggregation and compression graph-based techniques have been developed [172]. Sensor sampling schedulers have been designed using graph signal reconstruction techniques, so that the best subset of sensors is found so that the signal reconstruction error at the unobserved sensors is minimized, reducing the overall network energy consumption. Signal reconstruction techniques based on the GDFT have been used for the sensor sampling problem [70]. Recently, a graph sampling theory-based model has been developed for dynamic sensor placement, i.e., sensor are not static but can change positions [173].
5 Future Trends And Challenges
@@
In this section, we discuss some relevant challenges and trends of signal processing over graphs. However, we remark that applications reviewed such as virtual sensing, outlier detection, or missing value imputation are very active research fields. Table 6 depicts the challenges and future trends explained in the following sections.
Table 6: Summarization of the different challenges and future trends highlighted in the paper.
Challenges & Future Trends Graph Features
@@
(1) Transferability and Transfer Learning Graph-based models provide intradomain transferability and adaptation properties.
@@
(2) Federated and Distributed Learning Most of graph filters allow a distributed implementation.
@@
5.1 Generalization, Transferability, and Transfer Learning
@@
5.2 Federated and Distributed Learning
One of the relevant aspects of the GSP and graph filters is the fact that by explicitly using the graph shift matrix in their operations, which represents the relationship between nodes, they can be implemented in a distributed manner [185]. For instance, Xiao et al. [66] propose the graph Volterra series filter counterpart, which explicitly computes its output based on shifted versions of a graph signal and allows for a distributed approach. Other works have focused on improving the distributed application of graph filters by reducing the communication and computational resources required [36]. As in the case of graph filters, GNNs have been the focus of research in their distributed application, given the massive size of networks and data that are used to train such neural networks. Thus, it has been actively investigated how to distribute the training and applications of GNNs in clusters of machines trying to optimize the imbalance in communications and workload, in a parallel fashion [186, 187, 188]. For instance, Protogerou et al. [189] propose a distributed application of a GNN using multi-agents that work cooperatively to perform anomaly detection tasks.
@@ 5.3 Digital Twins
Digital twins are gaining popularity in the academic and industrial environment, and this technology is increasingly intended to be transferred to other fields, e.g., smart cities, e-health, precision agriculture, etc. A key component of digital twins is the sensing that complements an entity’s physical representation, giving feedback to improve an underlying physical model. Therefore, sensor networks play a key role in this paradigm and in particular the quality of the data they provide.
@@
Graph-based models, and more precisely GNNs, have been used to represent physical entities, therefore, producing digital twins. For instance, in the field of computer communications, GNNs have been used to predict network performance and behavior, being able to detect abnormal conditions [194, 195]. Ferriol-Galmés et al. [196] propose a GNN-based network digital twin to predict network performance and service level agreement metrics. Overall, we observe how IoT data can be fed to graph-based models, which can accurately represent physical entities and the relationship between the monitoring sites, to produce useful insights that can be later used for supporting decision-making processes, early warning, predictive maintenance, or anomaly detection. Furthermore, as seen in previous sections, graph-powered models can not only be effective in estimating digital twins but can also provide solutions to improve the operation of sensor networks that feed digital twins, e.g., through sensor identification and replacement [139]. In conclusion, digital twins can also be exploited to carry out applications such as anomaly or outlier detection as well as missing value imputation.
5.4 Quantum Graph-Based Models
@@
Recently with the rise of quantum computing, many quantum algorithms, or so-called quantum-inspired algorithms, have appeared that bring quantum mechanics fundamentals to machine learning algorithms to provide new properties, e.g., compact parameter representation, and faster training. Thus, much research has focused on developing the quantum analog of known techniques. This phenomenon has also reached graph-based models, particularly in the development of quantum GNNs. For instance, Verdon et al. [197] introduce the quantum GNNs to represent quantum processes that possess a graph structure. Quantum recurrent GNNs and quantum convolutional GNNs have also been proposed, however, these specific architectures are tailored to a given application, such as graph classification or others [198, 199]. Lately, quantum graph transformers have been proposed, which aim to translate and compute graph transformers in quantum computing primitives and processors [200]. In short, we observe how graph-based models such as GNN are translated into quantum-inspired models, i.e., models that bring ideas from quantum computing to improve their classical analog, and other models that aim to use a pure quantum approach with quantum processors. However, we see how a small number of applications have been tackled with quantum graph-based models, e.g., graph classification, and we believe that in the future more applications and tasks will benefit from quantum-inspired and pure-quantum approaches.
@@ 6 Conclusions
In this article, we have introduced some of the most relevant approaches in the field of data processing over graphs, i.e., graph signal processing, machine learning over graphs, and graph neural networks, for data quality in IoT monitoring sensor networks. Once the notions have been described, we have proposed a taxonomy of applications where these frameworks can be applied to data quality improvement in monitoring sensor networks, reviewing and unveiling new applications, e.g., air quality sensor networks, industrial sensor networks, or water management sensor networks. We have described some of the most novel tasks where signal processing and machine learning over graphs are making inroads, such as anomaly detection, missing value imputation, or even the creation of virtual sensors. In addition, we have classified the tasks into those approaches that work at the network level, at the edge level, and at the sensor level. We have discussed the main technologies and models, with emphasis on the models that have been developed more recently. Finally, in addition to the development of techniques for the various applications mentioned above, we have reviewed some of the most recent trends and challenges posed. We have reviewed the generalization and transfer learning capabilities that graph-based techniques provide, enabling future applications and solving data scarcity problems. We have emphasized the distributable nature and the possible federated learning use of such models. We have also presented the use of GNNs for the development of digital twins in some fields, e.g., computer communication, and we have highlighted their potential for monitoring sensor networks. Finally, we have introduced challenging future fields, e.g., quantum computing alternatives to improve model processing and scalability.
@@ Acknowledgments
This work is supported by Grant PID2022-138155OB-I00 funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF A way of making Europe”, CDTI MIG-20221061, and by regional project 2021SGR-01059.

### v2

@@ 4.3.4 Clustering and Data Aggregation
In the context of data aggregation and compression graph-based techniques have been developed [181]. Sensor sampling schedulers have been designed using graph signal reconstruction techniques, so that the best subset of sensors is found to minimize the signal reconstruction error at the unmonitored locations, reducing the overall network energy consumption. Signal reconstruction techniques based on the GDFT have been used for the sensor sampling problem [77]. Recently, a graph sampling theory-based model has been developed for dynamic sensor placement, i.e., sensors are not static but can change positions [182].
5 Challenges, Limitations, and Emerging Trends
@@
In this section, we discuss some relevant challenges and limitations that have arisen, as well as identify emerging trends in signal processing over graphs. However, we remark that the applications reviewed, such as virtual sensing, outlier detection, or missing value imputation, are very active research fields.
5.1 Challenges and Limitations
In this section, we describe some of the limitations and challenges that graph-based models face in modeling and improving IoT network data quality.
Large-scale IoT monitoring networks: one of the main challenges is the modeling of large-scale sensor networks. In this scenario, a graph can have thousands of nodes, so not only can the training of data quality enhancement models be complex, but the construction of the graph itself can be challenging. In the realm of GSP, linear algebra features, such as vectorization or parallelization of operations, have been used to define operations such as GDFT in high-dimensional scenarios [183]. In this line, specialized optimization techniques, such as the alternating direction method of multipliers (ADMM) have been used to solve large-scale modeling GSP-based problems [156, 61]. Moreover, GNNs provide operations, such as graph sampling or pooling, that allow efficient learning in high-dimensional graphs. In this regard, two fields have emerged that allow for optimizing training in large-scale networks, such as distributed training and the generalization and transference of models to larger networks (section 5.2.1).
Data availability: is one of the main requirements for training complex data-driven models. This aspect is also relevant in scenarios where IoT platforms are modeled with little data, but similar IoT data is available in other environments. Therefore, several research lines have been developed that study the transference of pre-trained models and the generalization of trained models on small networks to larger networks [14] (section 5.2.1).
Heterogeneity and dynamic graphs: Another aspect to consider is the complexity and heterogeneity of IoT networks. Although we have focused on static IoT monitoring networks in this survey, more advanced networks, such as mobile networks, not only have a variable number of nodes (since a node can enter or leave the network at any time), but also the relationships between nodes can change. Therefore, mobile monitoring networks would require the use of time-varying graphs [56]. Another aspect that has attracted attention is the assumption of smoothness and homophily used in graph-based models, where connected nodes tend to be similar. This assumption can be compromised in networks where sensors that measure different phenomena coexist or where there are different types of nodes. Recent studies analyze the performance of graph-based models such as GNNs in applications under heterophily [184].
Security: the security of graph models has been one of the most overlooked topics. As previously observed, the robustness of GNNs can be compromised by adversarial attacks [14]. Therefore, different attacks have been analyzed and the robustness of GNNs and GSP models has been the subject of study during recent years [185, 186, 187]. This topic is covered in section 5.2.5. Another related aspect is data privacy and ownership, with federated learning approaches being explored to ensure privacy in IoT applications (section 5.2.2).
5.2 Emerging Trends
In this section, we review some emerging trends in graph-based models for IoT monitoring networks. We show some of the trends that arise in response to some of the challenges and limitations presented in the previous section. Figure 15 represents the different emerging trends reviewed. Table 6 describes the graph features leveraged in the emerging trends explained in this section 5.2.
Figure 15: A number of emerging trends that take advantage of important features offered by graph-based models.
Table 6: Summarization of the different emerging trends highlighted in the paper.
Emerging Trends Graph Features
@@ 5.2 Emerging Trends
(1) Generalization and Transferability Graph-based models provide intradomain transferability and adaptation properties.
@@
(2) Distributed and Federated Learning Most of graph filters allow a distributed implementation.
@@
(5) Security Graph-based models need to provide robustness against cyberattacks.
5.2.1 Generalization, Transferability, and Transfer Learning
@@
5.2.2 Distributed and Federated Learning
One of the relevant aspects of the GSP and graph filters is the fact that by explicitly using the graph shift matrix in their operations, which represents the relationship between nodes, they can be implemented in a distributed manner [199]. For instance, Xiao et al. [72] propose the graph Volterra series filter counterpart, which explicitly computes its output based on shifted versions of a graph signal and allows for a distributed approach. Other works have focused on improving the distributed application of graph filters by reducing the communication and computational resources required [40]. The training and inference of distributed models in an IoT environment imply the adoption of an edge computing paradigm that needs to be aware of the resources available in the IoT nodes [200]. As in the case of graph filters, GNNs have been the focus of research in their distributed application, given the massive size of networks and data that are used to train such neural networks. Thus, it has been actively investigated how to distribute the training and applications of GNNs in clusters of machines trying to optimize the imbalance in communications and workload, in a parallel fashion [201, 202, 203]. For instance, Protogerou et al. [204] propose a distributed application of a GNN using multi-agents that work cooperatively to perform anomaly detection tasks.
@@ 5.2.3 Digital Twins
Digital twins are gaining popularity in the academic and industrial environment, and this technology is increasingly intended to be transferred to other fields, e.g., smart cities, e-health, precision agriculture, etc [209]. A key component of digital twins is sensing, which complements the physical representation of an entity and provides feedback to improve an underlying physical model. Therefore, sensor networks play a key role in this paradigm, especially the quality of the data they provide.
@@
Graph-based models, and more precisely GNNs, have been used to represent physical entities, therefore, producing digital twins. For instance, in the field of computer communications, GNNs have been used to predict network performance and behavior, being able to detect abnormal conditions [210, 211]. Ferriol-Galmés et al. [212] propose a GNN-based network digital twin to predict network performance and service level agreement metrics.
Applications have also been found in monitoring systems, where digital twins can help to detect different events and trigger the necessary actions. For instance, Roudbari et al. [213] use a GCRN in conjunction with a data-driven graph to predict water levels in a water monitoring network and feed a digital twin to assess potential flooding. Similarly, Sui et al. [214] use a graph-based digital twin to model a power system, in the so-called Internet of Energy (IoE), and apply a GCN to detect possible faults and estimate the stability of the system. In the realm of smart healthcare, GNNs have been used to estimate and forecast patients’ conditions [215].
In general, IoT data can be fed into graph-based models that accurately represent physical entities and relationships between monitoring sites to generate useful insights that can later be used to support decision-making processes, perform early warning, predictive maintenance, or anomaly detection. As noted in related publications, graph-based models can represent and provide real-time estimates for an entity or asset. These estimates can be used to detect possible events and failures, or even predict future conditions. Furthermore, as seen in previous sections, graph-powered models are not only effective in estimating digital twins but can also provide solutions to improve the operation of sensor networks that feed digital twins, e.g., through malfunctioning sensor identification and replacement [152]. Therefore, much recent work has focused on improving the data quality and operability of the sensor networks that feed these digital twins.
5.2.4 Quantum Graph-Based Models
@@
Recently with the rise of quantum computing, many quantum algorithms, or so-called quantum-inspired algorithms, have appeared that bring quantum mechanics fundamentals to ML algorithms to provide new properties, e.g., compact parameter representation, and faster training. Thus, much research has focused on developing the quantum analog of known techniques. This phenomenon has also reached graph-based models, particularly in the development of quantum GNNs. For instance, Verdon et al. [216] introduce the quantum GNNs to represent quantum processes that possess a graph structure. Quantum recurrent GNNs and quantum convolutional GNNs have also been proposed, however, these specific architectures are tailored to a given application, such as graph classification or others [217, 218]. Lately, quantum graph transformers have been proposed, which aim to translate and compute graph transformers in quantum computing primitives and processors [219]. In short, we observe how graph-based models such as GNN are translated into quantum-inspired models, i.e., models that bring ideas from quantum computing to improve their classical analog, and other models that aim to use a pure quantum approach with quantum processors. However, we see how a small number of applications have been tackled with quantum graph-based models, e.g., graph classification, and we believe that in the future more applications and tasks will benefit from quantum-inspired and pure-quantum approaches.
5.2.5 Security
Model security and robustness have been some of the most overlooked aspects in the development of graph-based models [14]. This is particularly interesting in an era where IoT data is being used in decision-making processes or, as seen above, in digital twin applications. Among the most known attacks are data poisoning, noise injection, or adversarial attacks on GNNs. In the case of GNNs, adversarial attacks affecting the robustness of the methods have been described. In particular, these attacks may include not only the modification of values but also the modification of the graph structure [220, 186, 185]. In terms of defense, just as attackers can use adversarial samples, adversarial training can be introduced into the GNN training methodology to provide robustness against possible attacks [220, 186]. Another defense method is to detect possible modifications of samples, nodes, or edges of the graph, and eliminate possible malicious elements. Similarly, in the context of GSP, GDFT-based techniques have been used to introduce structural adversarial samples, and various GSP methods, such as GDFT or smoothness, have been shown to be useful in detecting maliciously injected data [221, 187].
All in all, the models used must be robust to ensure the integrity of the IoT monitoring network data. In addition, detection models can be introduced to detect possible malicious data injection or modification.
@@ 6 Conclusions
This article has reviewed the use of graph-based models, based on frameworks such as GSP, ML over graphs, or GNNs, to improve data quality in IoT monitoring networks. The basics of the different graph-based frameworks that allow building the necessary models in the context of IoT have been presented. Furthermore, a taxonomy has been proposed to identify different applications that can be performed to improve data quality, such as missing value imputation, anomaly detection, or virtual sensing. All these applications can improve different dimensions of IoT data quality such as completeness, accuracy, or consistency. In addition, this taxonomy makes it possible to identify applications that practitioners can use, depending on their scenario, to meet some data quality requirements. The different applications have been classified according to the scope of the application, i.e., whether the objective is at the network level, at the edge or sensor relationship level, and at the node or sensor level. Once the different applications have been described, different limitations and challenges that may arise when using graph models have been identified. Challenges such as the use of models in large-scale IoT deployments, data availability and heterogeneity, and security. Finally, several emerging trends have been described, such as the development of digital twins using graph-based models or quantum graph models, as well as other areas that seek to provide solutions to the aforementioned limitations, e.g., the generalization and transfer of models, the distributed or federated training, or the security of graph-based models.
@@ Acknowledgments
This work is supported by Grant PID2022-138155OB-I00 funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF A way of making Europe”, CDTI MIG-20221061, regional project 2021SGR-01059, and by AGAUR 2023 CLIMA 0097.